# Improving Statistical Significance in Human Evaluation of Automatic Metrics via Soft Pairwise Accuracy

### v1

@@ Abstract
Selecting an automatic metric that best emulates human judgments is often non-trivial, because there is no clear definition of “best emulates.” A meta-metric is required to compare the human judgments to the automatic metric judgments, and metric rankings depend on the choice of meta-metric. We propose Soft Pairwise Accuracy (SPA), a new meta-metric that builds on Pairwise Accuracy (PA) but incorporates the statistical significance of both the human judgments and the metric judgments. SPA allows for more fine-grained comparisons between systems than a simplistic binary win/loss, and addresses a number of shortcomings with PA: it is more stable with respect to both the number of systems and segments used for evaluation, it mitigates the issue of metric ties due to quantization, and it produces more statistically significant results. SPA was selected as the official system-level metric for the 2024 WMT metric shared task.
@@ 1 Introduction
Automatic metrics are crucial because researchers and practitioners in NLP typically can’t always afford the high cost and latency of high-quality human evaluations. Despite their shortcomings, metrics like Word Error Rate (WER) and BLEU Papineni et al. (2002)—in conjunction with carefully curated test sets—have been crucial for the field of NLP, as they have provided a yardstick to make continual progress over many decades in Automatic Speech Recognition (ASR) and Machine Translation (MT), respectively.
@@
Reliance on automatic metrics makes selecting a good automatic metric of paramount importance. Conceptually, an automatic metric should emulate human judgments. Selecting an automatic metric typically entails generating a set of human judgments for a wide variety of outputs from a large number of different systems, and selecting the automatic metric that produces scores most similar to the human judgments. But what does “most similar” even mean? To select the “most similar” metric we need a meta-metric to compare metric scores to human judgments. The choice of meta-metric has been hotly debated in the field of MT (see § 5).
@@
Somewhat recently, Kocmi et al. (2021) made a case for choosing metrics with the best pairwise rankings—i.e. Pairwise Accuracy (PA) as a meta-metric—as the typical use case for a metric is to directly compare two systems on a given test set. This has been adopted by the WMT Metrics shared task for the past several years (Freitag et al., 2021, 2022, 2023). However, standard best practice when using an automatic metric is to consider not only which system a metric prefers, but also whether or not that preference is statistically significant. Thus we argue that metrics should not only emulate the accuracy of human pairwise ranking, they should also emulate the confidence or statistical significance of the human judgments.
@@
To this end, we propose Soft Pairwise Accuracy (SPA), which considers statistical significance of both the metric in question and the human judgments when evaluating the extent to which the metric agrees with the human judgments. We show that SPA, as its name implies, can be viewed as a soft (i.e. non-quantized) version of PA, and present analysis that demonstrates SPA has several distinct advantages over PA. In particular, we find it is more stable with respect to the number of segments and systems being scored, it does not produce a large number of ties between metrics, and it produces substantially more statistically significant comparisons between metrics. As a result of our recommendations, SPA is the official system-level metric for the 2024 WMT metric shared task.
@@ 2 Method
where $N$ is the number of systems, $p_{ij}^{h}$ is the $p$-value for hypothesis that system $i$ is better than system $j$ given the human judgments, and $p_{ij}^{m}$ is the $p$-value for hypothesis that MT system $i$ is better than system $j$ given the metric scores, and the first term is normalizing by the total number of pairs: (${{N\choose 2}}^{-1}=\frac{2}{N(N{-}1)}$. The $p$-value formulation will be defined in § 3.2, but for now, intuitively it can be thought of as an approximation to the confidence of the human raters and automatic metric. In particular, $p_{ij}\approx 0$ means the raters or metric are extremely confident that system $i$ is better than system $j$, $p_{ij}\approx 1$ means that the raters or metric are extremely confident that system $j$ is better than system $i$, and $p_{ij}\approx 0.5$ means that the raters or metric believe the two systems to be indistinguishable in terms of performance.
Figure 1: Illustration of the individual components used to calculate both pairwise confidence error and pairwise accuracy for the Prism metric (Thompson and Post, 2020a, b) on the WMT 2023 English-German language pair. Each box represents a comparison between two systems, systems $i$ and $j$. MT systems are sorted by average human judgment score for easier interpretation. The right column is one minus the absolute difference between the human preference for systems $i$ over system $j$ (left column) and the metric preference for system $i$ over system $j$ (middle column). In PA (top row), human and metric preferences are quantized to 0 and 1, and PA is thus an average of binary terms. In SPA (bottom row), human and metric preferences range from 0 to 1, and as a result SPA is an average of values ranging from 0 to 1. SPA can be viewed as a "soft" extension to pairwise accuracy that incorporates both human judgment and metric uncertainty, allowing for partial credit.
@@ 2.1 Relationship to Pairwise Accuracy
where $\,a_{ij}^{m}$ is 1 when the metric and human judgments agree and 0 otherwise.
A $p$-value $p_{ij}$ will be less than 0.5 when the human raters (or automatic metric) prefer system $i$ over system $j$, and greater than 0.5 when the human raters (or automatic metric) prefer system $j$ over system $i$.
This allows us to define PA in terms of quantized $p$-values:
@@
Where rounding is denoted as:
@@
Comparing Equation 1 and Equation 3, highlights the fact that SPA can be viewed as a ‘soft’ extension to pairwise accuracy that incorporates uncertainty in both the human and metric scores. A visualization of this is provided in Figure 1. We hypothesize that this rounding/quantization is effectively acting as additive random noise on top of the underlying SPA meta-metric.
In cases where both the MT metric and the human evaluation both have high statistical significance (regardless of whether the metric agrees with the human judgments or not, i.e. $p_{ij}^{m}\approx 0$ or $p_{ij}^{m}\approx 1$), the contribution of that system pair to SPA and PA is approximately identical. However, there are two important cases where our meta-metric differs from pairwise accuracy:
@@
    The human evaluation has high significance (e.g. $p_{ij}^{h}\approx 0$ or $p_{ij}^{h}\approx 1$), but the metric has low statistical significance (e.g. $p_{ij}^{m}\approx 0.5$): Even if the metric happens to choose the correct winner, we partially penalize the metric for not having high statistical significance.
@@
    The human evaluation finds the systems are approximately tied (e.g. $p_{ij}^{h}\approx 0.5$): In this case, we partially penalize the metric if has high statistical significance (e.g. $p_{ij}^{m}\approx 0$ or $p_{ij}^{m}\approx 1$) even if it happens to pick the same winner as the human evaluation, and to get full credit the metric must match the human evaluation statistical significance (e.g. $p_{ij}^{m}\approx p_{ij}^{m}\approx 0.5$)
@@ 2.2 Addressing Metric Ties in PA
The fact that PA considers only binary wins/losses (e.g. the rounding in Equation 3) results in an interesting shortcoming in PA. There are (${{N\choose 2}}$ pairs of $N$ systems, and thus (${{N\choose 2}}+1$ distinct values that PA can take on ($0/{N\choose 2}$, $1/{N\choose 2}$, …, (${N\choose 2}/{N\choose 2}$). By way of example, WMT-22 en-de, with $N=14$ systems, (${{N\choose 2}}+1=92$.
However, metrics tend to perform better than a random baseline, so only the upper half of the range is actually useful (e.g. this leaves 46 distinct values for $N=14$ systems). In practice, most metrics perform substantially better than random, and we tend to compare many (e.g. >10) metrics. In practice, we find that this results in PA reporting the same score for several sets of metrics (see § 4.3).
By removing this rounding/quantization, SPA has no such issues.
@@ 3.1 Data
We conduct experiments on the data from the WMT Metrics shared tasks, years 2022 and 2023 Freitag et al. (2022, 2023). In particular, we use the primary language pairs where MQM judgments were collected. We use the MT Metrics Eval V2 toolkit [^1] https://github.com/google-research/mt-metrics-eval to retrieve official shared task scores.
We make the fairly arbitrary decision to compare all metrics, including non-primary metrics but excluding QE metrics (i.e. reference-free metrics) which provide segment-level scores.
In order to compute the statistical significance of comparisons between metrics metrics, we make the simplifying assumption that all system-level metrics are the average of their segment-level metric. This is not true for some metrics, including BLEU Papineni et al. (2002) and chrF Popović (2015). While it would be possible to re-compute BLEU and chrF for each subset, we average for simplicity. Our results are self consistent (i.e. we use the segment-level average in all cases) but will not accurately reflect the true performance of these metrics. To the best of our knowledge, this approximation is also made in recent WMT shared tasks.
3.2 $p$-value Computation
Figure 2: Final metric ranking stability when ablating the number of MT systems (and thus the number of total MQM judgments), measured as change in Pearson correlation coefficient (Pearson $r$) from the ranking computed on all MT systems. We find SPA to be more stable than PA in all cases.
Given a pair of translations of the same test set, by two different MT systems, each scored at the segment level by the same metric, we would like to know the statistical significance of the difference in the means of the segment-level scores of each MT system.
We use a permutation test Fisher (1935) to estimate the probability that our observed difference occurred randomly under the null hypothesis that the two translations have the same quality. A permutation test works by generating random splits of the data (ignoring the labels, i.e. which MT system that produced each translation) and computing the difference in metric score mean for these two sets of data. The $p$-value is computed by calculating the fraction of the time that this difference is greater than or equal to the mean difference in scores for the two translations of the test set.
Permutation tests are appealing because they don’t require any assumptions about the underlying distribution of the data. This fits our use case well because we do not want to assume anything about the distribution of segment-level scores of a metric, which can obviously change depending on the particular metric. [^2] Metric and human annotation distributions are both highly variable Lo et al. (2023b); Knowles and Lo (2024). Permutation tests instead have the assumption of exchangeability Pitman (1937); Draper et al. (1993); Good (2002)—that is, under the null hypothesis (e.g. the MT systems are equally good) the joint distribution of the observations is invariant under permutations of the data labels. To help ensure exchangeability, we perform permutations such that each split has exactly one translation of each test set sentence, commonly referred to as a paired permutation test Good (2013). We estimate each $p$-value from 1000 permutations.
@@ 3.3 $p$-value Speed Optimization
A naive implementation of the paired permutation test is not computationally prohibitive when computing $p$-values for all systems/metrics a single time, but it becomes a problem when we want to compute these values many times in order to estimate statistical significance of metric comparisons.
Experimentally, we find the main speed bottleneck to be generating the random permutations, so when estimating statistical significance of metric comparisons we cache a batch of permutations and use it for each pair of systems, on a per test-set basis. Additionally, by sharing permutations across system pairs, this allows us to pre-compute the contribution of each system to means of the random permutations, allowing computations to be linear instead of square in the number of MT systems. See our code [^3] https://github.com/thompsonb/mt-metrics-eval/blob/main/mt_metrics_eval/pairwise_paired_permutation_test.py for full implementation details. This gave us an >1000x speedup over the default implementation in Scipy Virtanen et al. (2020). [^4] Using the ‘permutation_test’ function from scipy, ‘permutation_type’ set to ‘samples’ and the ‘n_resamples’ set to 1000, each $p$-value takes around 40 milliseconds to compute on a laptop CPU. Our approach has no impact on a $p$-value for a single system-level comparison, but it does mean that the $p$-value for one pair of systems is no longer independent from the $p$-value for any other pair of systems. Given our use case (i.e., we are using these $p$-values as an approximate level of confidence for the system-level comparisons in the meta-metric formulation, as opposed to making any claims about the actual statistical significance of the system-level comparisons) we believe this lack of independence should be inconsequential.
@@ 4 Analysis
Meta-metric evaluation is challenging because there is no ground truth (i.e., we don’t know the true ranking of the metrics). Instead, we must make some assumptions about how an ideal metric will act. First, we study how sensitive the meta-metric results are when varying the number of MT systems and number of segments per MT system, with the assumption that lower sensitivity to the exact systems / segments used indicates a better meta-metric. Second, we examine whether PA indeed has the problem of ties that we hypothesized, and whether SPA fixes this issue. Finally, we test our hypothesis that the quantization/rounding in PA is effectively acting as additive random noise, and that SPA is effectively the same underlying meta-metric with the noise term removed. If this is true (and the magnitude of the noise does not dominate the underlying signal), we would expect SPA to produce a similar metric ranking to PA, but with increased statistical significance.
@@ 4.1 Ablation: Number of Systems
Each year, WMT and the associated metrics task collect and score many online and submitted MT systems. For an ideal meta-metric, the exact choice of MT systems would have little to no impact on the metric rankings. We perform an ablation on the number of MT systems being scored, keeping the number of annotations per system fixed. We then compute the correlation (as measured by Pearson’s $r$) between the meta-metric’s ranking of the ablations compared to that same meta-metric’s full ranking. This allows us to evaluate how sensitive the metric is to the exact selection of MT systems.

### v2

@@ Abstract
Selecting an automatic metric that best emulates human annotators is often non-trivial, because there is no clear definition of “best emulates.” A meta-metric is required to compare the human judgments to the automatic metric scores, and metric rankings depend on the choice of meta-metric. We propose Soft Pairwise Accuracy (SPA), a new meta-metric that builds on Pairwise Accuracy (PA) but incorporates the statistical significance of both the human judgments and the metric scores. We show that SPA is more stable than PA with respect to changes in the number of systems/segments used for evaluation. We also show that PA can only assign a small set of distinct output values to metrics, and this results in many metrics being artificially assigned the exact same PA score. We demonstrate that SPA fixes this issue. Finally, we show that SPA is more discriminative than PA, producing more statistically significant comparisons between metrics. SPA was selected as the official system-level metric for the 2024 WMT Metrics Shared Task.
@@ 1 Introduction
Automatic metrics are crucial because researchers and practitioners in NLP typically can’t afford the high cost and latency of high-quality human evaluations. Despite their shortcomings, metrics like word error rate and BLEU Papineni et al. (2002)—in conjunction with carefully curated test sets—have been crucial for the field of NLP, as they have provided a yardstick to make continual progress over many decades in automatic speech recognition and machine translation (MT), respectively.
Figure 1: Illustration of the individual components used to calculate both SPA and PA for the Prism metric (Thompson and Post, 2020a, b) on the WMT 2023 English-German language pair. Each box represents a comparison between two systems, systems $i$ and $j$. MT systems are sorted by average human judgment score for easier interpretation. The right column is one minus the absolute difference between the human preference for systems $i$ over system $j$ (left column) and the metric preference for system $i$ over system $j$ (middle column). In PA (top row), human and metric preferences are binarized to 0 and 1, and PA is thus an average of binary terms. In SPA (bottom row), human and metric preferences range from 0 to 1, and as a result SPA is an average of values ranging from 0 to 1. SPA can be viewed as a "soft" extension to pairwise accuracy that incorporates both human judgment and metric uncertainty, allowing for partial credit.
@@
Reliance on automatic metrics makes selecting a good automatic metric of paramount importance. Conceptually, an automatic metric should emulate human judgments. Selecting an automatic metric typically entails generating a set of human judgments for a wide variety of outputs from a large number of different systems, and selecting the automatic metric that produces scores most similar to the human judgments. But how do we quantify similarity? To select the metric which produces output most similar to human judgements, we need a meta-metric to compare metric scores and human judgments. Despite nearly two decades of research on MT meta-evaluation, the community has not reached a consensus on the choice of a meta-metric. Various meta-metrics have been introduced over the years to address problems with prior meta-metrics, while sometimes creating new problems or re-introducing old ones (see § 5).
@@
Recent works Mathur et al. (2020b); Kocmi et al. (2021) have argued that the primary application of a metric is to choose between two competing systems, therefore the best metric is the one which produces pairwise system rankings most similar to the pairwise system rankings produced by human judgements. This led to Pairwise Accuracy (PA) being adopted by the WMT Metrics shared task for the past several years (Freitag et al., 2021, 2022, 2023). However, this argument omits a key detail: standard best practice when comparing two systems with an automatic metric is to consider not only which system the metric prefers, but also whether or not that preference is statistically significant Koehn (2004). Thus we argue that metrics should emulate not only the accuracy of human pairwise ranking, but also the confidence or statistical significance of the human pairwise ranking.
@@
To this end, we propose Soft Pairwise Accuracy (SPA), a new meta-metric which takes into account statistical significance of both the metric scores and the human judgments when evaluating the extent to which the metric in question agrees with the human judgments. We show that soft pairwise accuracy, as its name implies, can be viewed as a soft (i.e. non-binarized) version of PA, and present analysis that demonstrates SPA has several distinct advantages over PA. First, we find SPA is more stable with respect to the exact choice of MT systems and segments used. Second, we show that due to the binarization in its formulation, PA can only assign a small set of distinct output values to metrics, and in practice this results in many metrics being artificially assigned the exact same PA score. We demonstrate that SPA fixes this issue. Finally, we argue that PA is effectively equivalent to SPA with added noise due to binarization. We show that removing this noise (i.e. switching to SPA) results in substantially more statistically significant comparisons between metrics, making SPA a more discriminative and therefore more useful meta-metric. Our findings resulted in SPA being selected as the official system-level meta-metric for the 2024 WMT Metrics Shared Task Freitag et al. (2024).
@@ 2 Method
where $N$ is the number of systems for which we have human judgements and metric scores, $p_{ij}^{h}$ is the $p$-value for hypothesis that system $i$ is better than system $j$ given the human judgments, and $p_{ij}^{m}$ is the $p$-value for hypothesis that system $i$ is better than system $j$ given the metric scores. The term (${{N\choose 2}}^{-1}=\frac{2}{N(N{-}1)}$ normalizes the summation by the total number of pairs of systems being compared.
For each pairwise system comparison, we use a permutation test Fisher (1935) to estimate statistical significance of the difference in the means of the segment-level scores from a particular metric (or the human judgements) for the two systems. We first randomly split the segment-level scores (ignoring the labels, i.e. which MT system produced each segment) into two parts and compute the difference in metric score mean. Repeating this process many times provides a set of mean differences we can reasonably expect under the null hypothesis that the two systems are of the same quality. We compute a one-tailed $p$-value by calculating the fraction of the time that the random splits produce differences greater than or equal to the mean difference we observe for the two systems.
Permutation tests are appealing because they don’t require any assumptions about the underlying distribution of the data. This fits our use case well because we cannot assume anything about the distribution of segment-level scores of a metric. [^1] Metric and human annotation distributions are both highly variable Lo et al. (2023b); Knowles and Lo (2024). Permutation tests instead have the assumption of exchangeability Pitman (1937); Draper et al. (1993); Good (2002)—that is, under the null hypothesis (in our case, that the two MT systems are of equal quality) the joint distribution of the observations is invariant under permutations of the data labels. To help ensure exchangeability, we perform permutations such that each split has exactly one translation of each test set sentence, commonly referred to as a paired permutation test Good (2013).
Here we present some concrete examples for the sake of intuition. Suppose a metric reports a $+10$ point difference between system $i$ and system $j$, and that the random permutations only produce a metric difference $\geq 10$ points $1\%$ of the time. Thus $p_{ij}=0.01$ and we conclude that the metric has high confidence that system $i$ is better than system $j$. Likewise, if the metric reports the systems have a $-10$ point difference, we might find that the random permutations produce a metric difference $\geq-10$ points $99\%$ of the time. Thus $p_{ij}=0.99$ and we conclude the metric has high confidence that system $i$ is worse than system $j$. If the systems have the same metric score, we would expect about half of the random permutations to produce a metric difference $\geq 0$ and thus $p_{ij}=0.5$, indicating the metric finds the two systems indistinguishable from each other.
@@ 2.1 Relationship to Pairwise Accuracy
where $\,a_{ij}^{m}$ is 1 when the metric scores and human judgments prefer the same system and 0 otherwise. PA is equivalent to the Kendall rank correlation coefficient Kendall (1938), modulo a linear scaling and shifting (see § 5.1).
A $p$-value $p_{ij}$ will be less than 0.5 when the human raters (or automatic metric) prefer system $i$ over system $j$, and greater than 0.5 when the human raters (or automatic metric) prefer system $j$ over system $i$. This allows us to define PA in terms of binarized $p$-values:
@@
Where binarization is denoted as:
@@
Comparing Equation 1 and Equation 3 illustrates that SPA can be viewed as a ‘soft’ extension to PA that incorporates uncertainty in both the human and metric scores. A visualization of this is provided in Figure 1.
In cases where both the MT metric and the human evaluation both have high statistical significance (regardless of whether the metric agrees with the human judgments or not, i.e. $p_{ij}^{m}\approx 0$ or $p_{ij}^{m}\approx 1$), the contribution of that system pair to SPA and PA is approximately identical. However, there are two important cases where our meta-metric differs from PA:
@@
    The human evaluation has high statistical significance (i.e. $p_{ij}^{h}\approx 0$ or $p_{ij}^{h}\approx 1$), but the metric has low statistical significance (i.e. $p_{ij}^{m}\approx 0.5$): Even if the metric happens to choose the correct winner, we partially penalize the metric for not having high statistical significance.
@@
    The human evaluation finds the systems are approximately tied (i.e. $p_{ij}^{h}\approx 0.5$): In this case, we partially penalize the metric if has high statistical significance (i.e. $p_{ij}^{m}\approx 0$ or $p_{ij}^{m}\approx 1$) even if it happens to pick the same winner as the human evaluation, and to get full credit the metric must match the human evaluation statistical significance (i.e. $p_{ij}^{m}\approx p_{ij}^{m}\approx 0.5$)
@@ 2.2 Addressing Metric Ties in PA
The fact that PA considers only binary wins/losses (i.e. the binarization in Equation 3) results in an interesting shortcoming in PA. There are (${{N\choose 2}}$ pairs of $N$ systems, and thus (${{N\choose 2}}+1$ distinct values that PA can take on ($0/{N\choose 2}$, $1/{N\choose 2}$, …, (${N\choose 2}/{N\choose 2}$). For example, in WMT 2022 En-De, there are $N=14$ MT systems and thus (${{N\choose 2}}+1=92$.
However, metrics tend to perform better than a random baseline, so only the upper half of the range is actually useful (e.g. this leaves 46 distinct values for $N=14$ systems). We find that this results in PA reporting the same scores for several sets of metrics (see § 4.3). By removing this binarization, SPA has no such issues.
@@ 3.1 Data
We conduct experiments on the data from the 2022 and 2023 WMT Metrics Shared Tasks Freitag et al. (2022, 2023). In particular, we use the primary language pairs where MQM judgments were collected. We use the MT Metrics Eval V2 toolkit [^2] https://github.com/google-research/mt-metrics-eval to retrieve official shared task scores.
We make the somewhat arbitrary decision to compare all metrics, including non-primary metrics but excluding QE metrics (i.e. reference-free metrics) which provide segment-level scores.
In order to compute the statistical significance of comparisons between metrics, we make the simplifying assumption that all system-level metrics are the average of their segment-level metric. This is not true for some metrics, including BLEU Papineni et al. (2002) and chrF Popović (2015). While it would be possible to re-compute BLEU and chrF for each subset, we average the sentence-level versions of these metrics for simplicity. To the best of our knowledge, this approach is also taken in recent WMT metrics shared tasks.
Figure 2: Final metric ranking stability when ablating the number of MT systems (and thus the number of total MQM judgments), measured as change in Pearson correlation coefficient (Pearson $r$) from the ranking computed on all MT systems. Values are averaged over 1000 random trials. We find SPA to be more stable than PA in all cases.
@@ 3.2 $p$-value Speed Optimization
We estimate each $p$-value from 1000 random permutations. [^3] Note that due to the randomness inherent in the $p$-value estimation process, the exact value of SPA can vary slightly from run to run. A naive implementation of the paired permutation test is not computationally prohibitive when computing $p$-values for all systems/metrics a single time, but it becomes problematic when we want to compute these values many times in order to estimate statistical significance of metric comparisons.
Experimentally, we find the main speed bottleneck to be generating the random permutations, so when estimating statistical significance of metric comparisons we cache a batch of permutations and use it for each pair of systems, on a per test-set basis. Additionally, by sharing permutations across system pairs, this allows us to pre-compute the contribution of each system to means of the random permutations, allowing computations to be linear instead of square in the number of systems. See our code [^4] https://github.com/thompsonb/mt-metrics-eval/blob/main/mt_metrics_eval/pairwise_paired_permutation_test.py for full implementation details. This results in a speedup of over 1000x compared to the implementation in Scipy Virtanen et al. (2020). [^5] Using the ‘permutation_test’ function from scipy, ‘permutation_type’ set to ‘samples’ and the ‘n_resamples’ set to 1000, each $p$-value takes around 40 milliseconds to compute on a laptop. Our speed optimization does not change the computation of a $p$-value for a single system-level comparison, but it does mean that the $p$-value for one pair of systems is no longer computed independent from the $p$-value for any other pair of systems. Given that we are using these $p$-values as an approximate level of confidence for the system-level comparisons in the SPA meta-metric formulation, as opposed to making any claims about the actual statistical significance of the system-level comparisons, we believe this lack of independence should be inconsequential.
@@ 4 Analysis
Meta-metric evaluation is challenging because there is no ground truth (i.e., we don’t know the true ranking of the metrics). Instead, we conduct analysis to compare SPA and PA. First, we study how sensitive the meta-metric results are when ablating the number of MT systems and number of segments per MT system, with the assumption that lower sensitivity to the exact systems/segments used indicates a better meta-metric. Second, we examine whether PA indeed has the problem of ties that we hypothesized in § 2.2, and whether SPA fixes this issue. Finally, we test our hypothesis that the binarization in PA is effectively acting as additive random noise, and that SPA is effectively the same underlying meta-metric with the noise term removed.
@@ 4.1 Ablation: Number of Systems
Each year, WMT and the associated metrics task collect and score many online and submitted MT systems. For an ideal meta-metric, the exact choice of MT systems would have minimal impact on the metric rankings. We perform an ablation on the number of MT systems being scored, keeping the number of annotations per system fixed. We then compute the correlation (as measured by Pearson’s $r$) between the meta-metric’s ranking of the ablations compared to that same meta-metric’s full ranking. This allows us to evaluate how sensitive the metric is to the exact selection of MT systems.